The nursing instructor is discussing asthma and its various treatments with the students. The instructor tasks the students with preparing a patient teaching brochure about the use of the leukotriene receptor antagonists (LTRA) montelukast (Singulair®) in the treatment of asthma.

1. What information will the students include in the brochure regarding the adverse effects of montelukast?

2. What information will the students include in the brochure regarding drug interactions that can occur with montelukast?

3. What are three other patient teaching points that the students should include in the brochure regarding this medication?

4. The physician ordered 1g of Cefuroxime IV. You have on hand 125mg/5ml vial of Cefuroxime. How many mls of Cefuroxime will you administer?

5. Physician order reads: administer Cefuroxime IV over one hour in a 100 ml bag. The drop factor of tubing is 60 gtts/ml. What is the drip rate?
Answer #1

ANSWER.

1) Adverse effects of Montelukast.

           * Headache.

           * Abdominal pain.

           * Cough.

           * Diarrhea.

           * Weakness.

           * Earache..

           * Rhinitis.

           * Sore throat.

2) Drug interactions occur with Montelukast.

        Drugs which make interactions with montelukast includes

              * Losartan.

              * Amiodarone.

              * Rifampicin.

              * Dexamethasone.

              * Trimethoprim.

              * Zafirlukast.

              * Fluconazole.

              * Phenytoin.

              * Amodiaquine.

3) Three other teaching points regarding Montelukast.

           * Indications of montelukast.

               it is used to treat

                  * Allergic rhinitis.

                * Asthma.

                  * Bronchospasm.

        * Routes of administration.

                * Oral route (Tablets).

         * Containdications for montelukast.

                 * Depression.

                 * Hepatitis.

                 * Severe asthma.

                 * Jaundice.

                 * Pregnancy.

                 * Phenylketonuria.

4) Amount(ml) of Cefuroxime need to be administered.

         Formula = Desired dose/Dose in hand * Quantity.

            Here

               Desired dose = 1 gm.

         Convert gm to mg for calculation.

                   1gm = 1000mg.

    So

           Desired dose(D) = 1000 mg.

           Dose in hand(H) = 125 mg.

                Quantity(Q) = 5 ml.

As per formula

     Amount(ml) of drug required = D/H * Q

                                                   = 1000/125 * 5

                                                   = 8 * 5

                                                   = 40 ml.

5) Calculation of drip rate.

       Formula

           Drip rate(gtt/min) = Volume(ml) / Time(min) * Drop factor.

        Here

            Volume(ml) = 100 ml.

             Time(min) = 60 min.(1 hr)

             Drop factor = 60 gtts/ml.

     As per formula

             Drip rate(gtt/min) = 100/60 * 60

                                         = 100 gtt/min.